# Long-Context Autoregressive Video Modeling with Next-Frame Prediction

### v1

@@ Abstract
Long-context autoregressive modeling has significantly advanced language generation, but video generation still struggles to fully utilize extended temporal contexts. To investigate long-context video modeling, we introduce Frame AutoRegressive (FAR), a strong baseline for video autoregressive modeling. Just as language models learn causal dependencies between tokens (i.e., Token AR), FAR models temporal causal dependencies between continuous frames, achieving better convergence than Token AR and video diffusion transformers. Building on FAR, we observe that long-context vision modeling faces challenges due to visual redundancy. Existing RoPE lacks effective temporal decay for remote context and fails to extrapolate well to long video sequences. Additionally, training on long videos is computationally expensive, as vision tokens grow much faster than language tokens. To tackle these issues, we propose balancing locality and long-range dependency. We introduce FlexRoPE, an test-time technique that adds flexible temporal decay to RoPE, enabling extrapolation to 16$\times$ longer vision contexts. Furthermore, we propose long short-term context modeling, where a high-resolution short-term context window ensures fine-grained temporal consistency, while an unlimited long-term context window encodes long-range information using fewer tokens. With this approach, we can train on long video sequences with a manageable token context length. We demonstrate that FAR achieves state-of-the-art performance in both short- and long-video generation, providing a simple yet effective baseline for video autoregressive modeling.
@@ 1 Introduction
Building on FAR, we investigate long-context video modeling and explore two common settings, similar to those in language modeling: (1) test-time lengthy extrapolation and (2) long-sequence training. Unlike language modeling, long-context video modeling suffers from visual redundancy. On one hand, RoPE [51] exhibits weak temporal decay, leading to the accumulation of redundant visual context and degrading test-time extrapolation performance. On the other hand, training on long videos is computationally expensive, as vision tokens grow significantly faster than language tokens.
To address these challenges, we propose to balance the locality and long-range dependency. For test-time temporal extrapolation, we introduce FlexRoPE, which incorporates a controllable temporal decay into RoPE using a linear bias. FlexRoPE is applied only at test time and remains compatible with models trained using RoPE. It enhances temporal locality and effectively reduces redundancy from distant contexts while still allowing the learned RoPE to model long-range dependencies. For long-video training, we introduce long short-term context modeling. Specifically, we maintain a high-resolution short-term context window to ensure fine-grained temporal consistency, while using a unlimited long-term context window with aggressive patchification to reduce redundant context tokens. This strategy enables efficient training on long video sequences with a manageable token context length.
@@
    Building on FAR, we introduce FlexRoPE to enable 16$\times$ test-time temporal extrapolation, and long short-term context modeling for efficient long-video training.
@@
    FAR achieves state-of-the-art performance in both short- and long-video generation.
@@ 2.2 Long-Context Language Modeling
Test-time Lengthy Extrapolation. Lengthy extrapolation is an appealing characteristic of autoregressive models, allowing them to be trained on short sequences while performing inference on longer ones. However, extrapolation performance primarily depends on the characteristics of the position embedding. Two common relative position embeddings that support extrapolation are RoPE [51] and ALiBi [48]. RoPE encodes relative distance through dot-product operations, while ALiBi achieves this using attention bias. Subsequent studies [47, 6] have further advanced RoPE to enhance extrapolation performance. In this work, we introduce FlexRoPE and demonstrate its superior performance compared to RoPE and ALiBi in temporal extrapolation for video modeling.
Table 1: Model Variants of FAR. We follow the model size configurations of DiT [46] and SiT [38].
@@
Long Sequence Training. A straightforward approach to improving long-context modeling performance is to directly train the model on longer sequences. Recent work in language modeling has explored efficient long-sequence fine-tuning with position interpolation [12, 13]. However, vision tokens scale much faster than language tokens as context increases. To address this, we introduce long short-term context modeling to eliminate visual redundancy in long-video training.
2.3 Long-Context Video Modeling
Recent advancements in video generation models have enabled their use as interactive world simulators [53, 9, 45], which require the ability to exploit long-range context and memorize the observed environment. However, existing video diffusion transformers lack effective mechanism to utilize long-range context. Although early work [60] has explored long-context video prediction, it has been limited in visual quality and long-range consistency. In this work, we introduce FAR, a scalable framework for both short- and long-context autoregressive video modeling.
@@ 4 FAR
4.1 Framework Overview
Architecture. As shown in Fig. 2 (a), FAR is built upon the diffusion transformer [38, 46]. We adopt the model configuration of DiT [46] and Latte [39], as listed in Tab. 1. The key architectural difference between FAR and video diffusion transformers (e.g., Latte [39]) lies in the attention mechanism. As shown in Fig. 3(a), for each frame, we apply causal attention at the frame level while maintaining full attention within each frame. We adopt this causal spatiotemporal attention for all layers, instead of the interleaved spatial and temporal attention used in Latte. In FAR, image generation and image-conditioned video generation are jointly learned thanks to the causal mask, whereas video diffusion transformer [39] requires additional image-video co-training.
Figure 6: Visualization and Comparison of Various Temporal Position Embeddings. The proposed FlexRoPE incorporates a linear bias to induce controllable temporal decay at test time, enhancing extrapolation performance as context increases. In contrast, other methods degrade when the inference context exceeds the training window.
Basic Training Pipeline. The training pipeline of FAR is illustrated in Fig. 2 (a). Given a video sequence $\mathbf{X}$, we first employ a pretrained VAE to compress it into the latent space $\mathbf{Z}\in\mathbb{R}^{T\times H\times W}$, where $T$, $H$, and $W$ denote the number of frames, height, and width of the latent features, respectively. Note that although we primarily adopt an image VAE in this work, FAR can also be trained with a video VAE since our autoregressive unit is the latent frame. Following diffusion forcing [10], we independently sample a timestep for each frame. We then interpolate between the clean latent and the sampled noise using Eq. (1) and apply the frame-wise flow matching objective in Eq. (3) for learning. The key difference between FAR and image flow matching lies in that we adopt causal spatiotemporal attention, allowing each frame to access previous context frames during denoising.
@@ 4.2 Short-Video Modeling
Training-Inference Gap in Observed Context. As a hybrid AR-diffusion model, FAR also encounters a training-inference gap in the observed context. As illustrated in Fig. 2(a), each clean latent is fused with sampled noise for the flow matching objective, as defined in Eq. (1). Consequently, later frames can only access the noised version of previous frames during training. However, during inference, this leads to a distribution shift when clean context frames is used.
@@ 4.3 Long-Context Video Modeling
In this section, we discuss the challenge of long-context video modeling. We focus on two practical long-context settings, similar to those in language models: test-time temporal extrapolation (Sec. 4.3.1) and long-video training (Sec. 4.3.2).
4.3.1 Test-Time Temporal Extrapolation
Weak Temporal Decay. An appealing characteristic of autoregressive models is their potential to be trained on short sequences while being tested on long sequences, enabling lengthy extrapolation at test time. This capability relies on effective positional embedding. Following 3D-RoPE [61] for video data, the spatial and temporal dimensions (i.e., height, width, and frame) are treated as independent 1D-RoPE embeddings. Consequently, we keep RoPE positional embedding for height and width unchanged while focusing only on the temporal position embedding. In this study, we examine RoPE [51] and ALiBi [48], two common positional embedding methods for capturing temporal relative distances. From the visualization in Fig. 6(a), RoPE does not impose sufficient temporal decay, leading to accumulated redundant visual context. Similarly, ALiBi exhibits this issue in parts of attention heads with small slopes.
Figure 7: Relation between Token Context Length and Vision Context Length. With the proposed long short-term context modeling, the token context length scales more slowly as the vision context length increases compared to uniform context modeling.
To evaluate extrapolation performance, we gradually increase the number of context frames at test time and measure the resulting improvement in predictions. From Fig. 6(b), position interpolation of RoPE performs slightly better than position extrapolation. Additionally, we retrain the model using ALiBi as temporal position embedding. ALiBi applies a linear decay based on temporal relative distance, with different decay rates assigned to each attention head. Our results suggest that ALiBi achieves slightly better extrapolation than RoPE due to its explicit temporal decay mechanism. However, all solutions exhibit performance degradation as vision context increases. Therefore, we aim to develop a more effective temporal position embedding method to improve temporal extrapolation.
Our Solution: FlexRoPE. To address this problem, we propose FlexRoPE, which explicitly controls temporal decay to suppress redundant visual context while still allowing the model to capture long-range dependencies using RoPE. The FlexRoPE is defined as:
$\text{Attention}(q_{i},k_{j})=\text{Softmax}\left(\underbrace{\overbrace{RoPE(q_{i},k_{j})}^{\text{used in training}}-\lambda\cdot|i-j|}_{\text{FlexRoPE, used in inference}}\right),$ (5)
where the temporal decay is flexibly controlled by the slope $\lambda$, and $i$ and $j$ represent the temporal indices of the frame. We visualize FlexRoPE with $\lambda=0.2$ in Fig. 6(a). FlexRoPE is inference-compatible with models trained using RoPE since it does not modify the dot-product computation but instead compensates for RoPE’s temporal decay.
We compare FlexRoPE with RoPE in Fig. 6(b). FlexRoPE effectively balances locality and long-range correspondence, leading to improved performance as the context frame increases, whereas RoPE interpolation and extrapolation exhibit poorer extrapolation performance.
4.3.2 Long-Video Training
@@
Token Redundancy in Long Video. Visual data contains spatial redundancy, causing vision tokens to expand much faster than language tokens as context increases. For example, a video sequence of 128 frames requires more than 8K tokens, with 64 tokens per frame, as illustrated in Fig. 7. As a result, training and inference on long videos become computationally inefficient.
@@
Our Solution: Long Short-Term Context Modeling. To address the token redundancy in video, we introduce long short-term context modeling, which exploits the spatial and temporal locality in video data. As illustrated in Fig. 2(b), we maintain a high-resolution short-term context window to learn fine-grained temporal consistency and a low-resolution long-term context window, where we adopt aggressive patchification to reduce the number of context tokens. During training, given that the data has a maximum sequence length of $m$ frames, we fix the short-term context window to $n$ frames and randomly sample the long-term context frames from the range $[0,m-n]$. The attention mask with long short-term context modeling is shown in Fig. 3(b), where the long-term context uses fewer tokens. As demonstrated in Fig. 7, this strategy ensures that increasing the vision context length maintains a manageable token context length. To prevent interference between long-term and short-term contexts, we adopt separate projection layers for each context, inspired by MM-DiT [17]. This approach results in a slightly larger parameter size during long-video training, as shown in Tab. 1.
@@
Table 3: Quantitative Comparison on Short Video Prediction. We follow the evaluation setup of MCVD [54] and ExtDM [63], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
@@
Table 4: Quantitative Comparison on Long-Context Video Prediction. We follow the evaluation setup of TECO [60], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
@@
5 Experiment
5.1 Implementation Details
We follow the DiT’s structure [46] to implement FAR. To compress video latents, we train a series of image DC-AE [11] on the corresponding dataset, resulting in 64 tokens per frame. All models are trained from scratch without image pretraining. We provide detailed settings in Tab. 7.
5.2 Quantitative Comparison
5.2.1 Video Generation
Dataset and Evaluation Setting. We benchmark both unconditional and conditional video generation on the UCF-101 dataset [50], which consists of approximately 13,000 videos. Following Latte [39], we use the entire dataset for training. For evaluation, we randomly sample 2,048 videos to compute FVD [52] against the ground-truth videos. For conditional video generation, we set the guidance scale to 2.0 during inference.
Main Results. From the results listed in Tab. 2, we achieve state-of-the-art performance in both unconditional and conditional video generation. Specifically, Latte [39] is based on video diffusion transformer, while OmniTokenizer [56] is based on Token AR. Our method significantly outperforms both. Furthermore, compared to recent frame-autoregressive models [30, 65], which require twice the training cost, FAR achieves superior performance without any additional training cost.
@@ 5.2.2 Short-Video Prediction
Main Results. We summarize the results in Tab. 3. Unlike previous works such as MCVD [54] and ExtDM [63], which introduce complex multi-scale fusion strategies and optical flow, FAR achieves superior results on both datasets without requiring additional design.
@@ 5.2.3 Long-Video Prediction
Main Results. We summarize the results in Tab. 4. The previous work, TECO [60], adopts aggressive downscaling for all frames to reduce tokens for temporal modeling, creating a trade-off between training efficiency and prediction accuracy. In contrast, FAR employs long short-term context modeling, effectively achieving the lowest prediction error (i.e., LPIPS) without prohibitive computation cost.

### v2

@@ Abstract
Long-context autoregressive modeling has significantly advanced language generation, but video generation still struggles to fully utilize extended temporal contexts. To investigate long-context video modeling, we introduce Frame AutoRegressive (FAR), a strong baseline for video autoregressive modeling. Just as language models learn causal dependencies between tokens (i.e., Token AR), FAR models temporal causal dependencies between continuous frames, achieving better convergence than Token AR and video diffusion transformers. Building on FAR, we observe that long-context video modeling faces challenges due to visual redundancy. Training on long videos is computationally expensive, as vision tokens grow much faster than language tokens. To tackle this issue, we propose balancing locality and long-range dependency through long short-term context modeling. A high-resolution short-term context window ensures fine-grained temporal consistency, while an unlimited long-term context window encodes long-range information using fewer tokens. With this approach, we can train on long video sequences with a manageable token context length, thereby significantly reducing training time and memory usage. Furthermore, we propose a multi-level KV cache designed to support the long short-term context modeling, which accelerating inference on long video sequences. We demonstrate that FAR achieves state-of-the-art performance in both short- and long-video generation, providing a simple yet effective baseline for video autoregressive modeling. The code is released at https://github.com/showlab/FAR.
Index Terms:
Video Generation, Autoregressive Video Modeling, Diffusion Model.
@@ 1 Introduction
Building on FAR, we investigate how to train on long video sequences to enable long-context video modeling. Unlike language modeling, long-context video modeling suffers from visual redundancy, as vision tokens grow significantly faster than language tokens when context increases. To address this challenge, we propose to balance the locality and long-range dependency with long short-term context modeling. Specifically, we maintain a high-resolution short-term context window to ensure fine-grained temporal consistency, while using a unlimited long-term context window with aggressive patchification to reduce redundant context tokens. This strategy enables efficient training on long video sequences with a manageable token context length. To speed up the proposed long short-term context modeling during inference, we introduce a multi-level KV cache to enhance inference efficiency.
@@
    Building on FAR, we introduce long short-term context modeling for efficient long-video training, as well as multi-level KV cache for faster inference.
@@
    FAR enables significantly lower training costs on long videos and achieves state-of-the-art performance in both short- and long-video generation.
@@ 2.2 Long-Context Language Modeling
A straightforward approach to improving long-context ability in language modeling is to directly fine-tune the model on longer sequences. Recent work [33, 34] has explored efficient long-sequence fine-tuning with position interpolation. However, training on long videos leads to prohibitive computational costs, as vision tokens scale much faster than language tokens with increasing context length. To address this issue, we introduce long short-term context modeling to reduce visual redundancy in long-video training.
2.3 Long-Context Video Modeling
Recent advancements in video generation models have enabled their use as interactive world simulators [35, 36, 37], which require the ability to exploit long-range context and memorize the observed environment. However, existing video diffusion transformers lack effective mechanism to utilize long-range context. Although early work [38] has explored long-context video prediction, it has been limited in visual quality and long-range consistency. In this work, we introduce FAR, a efficient framework for both short- and long-context autoregressive video modeling.
TABLE I: Model Variants of FAR. We follow the model size configurations of DiT [39] and SiT [40].
@@ 4 FAR
4.1 Framework Overview
Architecture. As shown in Fig. 2 (a), FAR is built upon the diffusion transformer [40, 39]. We adopt the model configuration of DiT [39] and Latte [44], as listed in Table. I. The key architectural difference between FAR and video diffusion transformers (e.g., Latte [44]) lies in the attention mechanism. As shown in Fig. 3(a), for each frame, we apply causal attention at the frame level while maintaining full attention within each frame. We adopt this causal spatiotemporal attention for all layers, instead of the interleaved spatial and temporal attention used in Latte. In FAR, image generation and image-conditioned video generation are jointly learned thanks to the causal mask, whereas video diffusion transformer [44] requires additional image-video co-training.
Basic Training Pipeline. The training pipeline of FAR is illustrated in Fig. 2 (a). Given a video sequence $\mathbf{X}$, we first employ a pretrained VAE to compress it into the latent space $\mathbf{Z}\in\mathbb{R}^{T\times H\times W}$, where $T$, $H$, and $W$ denote the number of frames, height, and width of the latent features, respectively. Note that although we primarily adopt an image VAE in this work, FAR can also be trained with a video VAE since our autoregressive unit is the latent frame. Following diffusion forcing [17], we independently sample a timestep for each frame. We then interpolate between the clean latent and the sampled noise using Eq. Eq. 1 and apply the frame-wise flow matching objective in Eq. Eq. 3 for learning. The key difference between FAR and image flow matching lies in that we adopt causal spatiotemporal attention, allowing each frame to access previous context frames during denoising.
@@ 4.1 Framework Overview
Figure 6: KV Cache for Short-Video Modeling in FAR. We additionally add a caching step to encode current decoded frame into the KV cache for autoregressive generation.
Figure 7: Relation between Token Context Length and Vision Context Length. With the proposed long short-term context modeling, the token context length scales more slowly with increasing vision context length compared to uniform context modeling. When training on long videos, the reduced number of tokens leads to significantly lower training costs and memory usage.
@@ 4.2 Short-Video Modeling
Training-Inference Gap in Observed Context. As a hybrid AR-diffusion model, FAR also encounters a training-inference gap in the observed context. As illustrated in Fig. 2(a), each clean latent is fused with sampled noise for the flow matching objective, as defined in Eq. Eq. 1. Consequently, later frames can only access the noised version of previous frames during training. However, during inference, this leads to a distribution shift when clean context frames is used.
@@
Figure 8: Multi-Level KV Cache for Long-Context Video Modeling in FAR. When a frame leaves the short-term context window, we encode it to the L2 cache and re-encode the L1 cache in the window. We then use those encoded KV cache for decoding the current frame. Note that we divide the process into three steps for better illustration, though it can be merged into a single forward pass in implementation.
@@ 4.3 Long-Context Video Modeling
Token Redundancy in Long Video. Visual data contains spatial redundancy, causing vision tokens to expand much faster than language tokens as context increases. For example, a video sequence of 128 frames requires more than 8K tokens, as illustrated in Fig. 7(a). Consequently, training on long videos becomes computationally prohibitive, as shown in Fig. 7(b, c).
@@
Our Solution: Long Short-Term Context Modeling. To address the token redundancy in video, we introduce long short-term context modeling, which exploits the spatial and temporal locality in video data. As illustrated in Fig. 2(b), we maintain a high-resolution short-term context window to learn fine-grained temporal consistency and a low-resolution long-term context window, where we adopt aggressive patchification to reduce the number of context tokens. During training, given that the data has a maximum sequence length of $m$ frames, we fix the short-term context window to $n$ frames and randomly sample the long-term context frames from the range $[0,m-n]$. The attention mask with long short-term context modeling is shown in Fig. 3(b), where the long-term context uses fewer tokens. As demonstrated in Fig. 7(a), this strategy ensures that increasing the vision context length maintains a manageable token context length. With long short-term context modeling, we can reduce the cost and memory usage of the long-video training significantly, as shown in Fig. 7(b, c). To prevent interference between long-term and short-term contexts, we adopt separate projection layers for each context, inspired by MM-DiT [45]. This approach results in a slightly larger number of parameters, referred to as FAR-Long in Table. I.
TABLE II: Experimental Configurations of FAR. We follow the evaluation settings from Latte [44], MCVD [46], and TECO [38].
Hyperparameters Short-Video Generation Short-Video Prediction Long-Video Prediction
Cond. UCF-101 Uncond. UCF-101 BAIR UCF-101 Minecraft DMLab
Dataset Configuration
Resolution 256/128 256/128 64 64 128 64
Total Training Samples 13,320 13,320 43,264 9,624 194,051 39,375
Training Configuration
Batch Size 32 32 32 32 32 32
Latent Size 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47]) 8$\times$8 (DC-AE [47])
Training Sequence Length 16 16 32 16 300 300
LR $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
LR Schedule constant constant constant constant constant constant
Warmup Steps - - - - 10K 10K
Total Training Steps 400K 400K 200K 200K 1M 1M
Stochastic Clean Context 0.1 0.1 0.1 0.1 0.1 0.1
Short-Term Context Window 16 16 32 16 16 16
Long-Term Context Resolution - - - - 2$\times$2 2$\times$2
Evaluation Configuration
Samples 4$\times$2048 4$\times$2048 100$\times$256 100$\times$256 4$\times$256 4$\times$256
Guidance Scale 2.0 - - - 1.5 1.5
Reference Work Latte [44] Latte [44] MCVD [46] MCVD [46] TECO [38] TECO [38]
4.4 Inference-Time KV Cache
KV Cache for Short-Video Modeling. Due to the autoregressive nature of FAR, we can leverage KV-Cache to accelerate inference. As illustrated in Fig. 6, for each frame, we first use the flow-matching schedule to decode it into the clean latent frame. We then introduce an additional caching step to encode the clean latent frame into the KV cache. As discussed in Sec. 4.2, we use timestep $t=-1$ to denote the clean context frame in the caching step. These KV caches are subsequently used for autoregressive decoding of subsequent frames.
Multi-Level KV Cache for Long-Video Modeling. In long-context video modeling, we employ long short-term context to reduce redundant visual tokens. To accommodate this, we introduce a multi-level KV cache. As illustrated in Fig. 8, the frames in long-term context window is encoded into L2 cache (4 tokens per frame), while the frames in short-term context window is encoded into L1 cache. When decoding current frame but exceed the short-term context window, the earliest frame in the short-term context window is moved to the long-term context window and encode it into the L2 cache. Since this modifies the cache state, we subsequently re-encode the L1 cache of the frames in the short-term context window. The encoded cache is then used to decode the current frame. Note that in practice, these three steps can be merged into a single forward pass for efficiency.
5 Experiment
5.1 Implementation Details
We follow the DiT’s structure [39] to implement FAR. To compress video latents, we train a series of image DC-AE [47] on the corresponding dataset, resulting in 64 tokens per frame. All models are trained from scratch without image pretraining. We provide detailed training hyperparameters and evaluation setting in Table. II.
5.2 Quantitative Comparison
5.2.1 Video Generation
Dataset and Evaluation Settings. We benchmark both unconditional and conditional video generation on the UCF-101 dataset [48], which consists of approximately 13,000 videos. Following Latte [44], we use the entire dataset for training. For evaluation, we randomly sample 2,048 videos to compute FVD [49] against the ground-truth videos. For conditional video generation, we set the guidance scale to 2.0 during inference.
Main Results. From the results listed in Table. III, we achieve state-of-the-art performance in both unconditional and conditional video generation. Specifically, Latte [44] is based on video diffusion transformer, while OmniTokenizer [50] is based on Token AR. Our method significantly outperforms both. Furthermore, compared to recent frame-autoregressive models [20, 21], which require twice the training cost, FAR achieves superior performance without any additional training cost.
@@ 5.2.1 Video Generation
TABLE IV: Quantitative Comparison on Short Video Prediction. We follow the evaluation setup of MCVD [52] and ExtDM [46], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
@@
TABLE V: Quantitative Comparison on Long-Context Video Prediction. We follow the evaluation setup of TECO [38], where $c$ denotes the number of context frames and $p$ denotes the number of predicted frames.
@@ 5.2.2 Short-Video Prediction
Main Results. We summarize the results in Table. IV. Unlike previous works such as MCVD [52] and ExtDM [46], which introduce complex multi-scale fusion strategies and optical flow, FAR achieves superior results on both datasets without requiring additional design.
@@ 5.2.3 Long-Video Prediction
Main Results. We summarize the results in Table. V. The previous work, TECO [38], adopts aggressive downscaling for all frames to reduce tokens for temporal modeling, creating a trade-off between training efficiency and prediction accuracy. In contrast, FAR employs long short-term context modeling, effectively achieving the lowest prediction error (i.e., LPIPS) without prohibitive computation cost.
@@ 5.3 Qualitative Comparison
Figure 9: Ablation Study of the KV Cache. FAR-Long with proposed multi-level KV cache achieves the best speedup on long videos.